Consider the problem: According to last year's report, a Filipino household spends an average of P500 per day for food. Suppose you recently took random samples of 30 households. You determined how much each households spent for food each day and the results revealed a mean of P450 and the standard deviation of P25.50. Using a 0.05 level of significance, can it be concluded that the average amount spent per day for food of a Filipino household has decreased? Assume normality over the population

1
Expert's answer
2022-05-22T17:50:01-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge500"

"H_1:\mu<500"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=29" degrees of freedom, and the critical value for a left-tailed test is "t_c = -1.699127."

The rejection region for this left-tailed test is "R = \{t:t<-1.699127\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{450-500}{25.50/\sqrt{30}}\approx-10.73968"

Since it is observed that "t=-10.73968<-1.699127=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed, "df=29" degrees of freedom, "t=-10.73968," is "p= 0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 500, at the "\alpha = 0.05" significance level.